Assume that the price of a new model of car in is ₹4,60,000 and 90,000 cars are sold at this price in a year. If the price elasticity of demand of new car is 1.7, what will be the effect on annual sales when the average price of new model declines to ₹4,15,000?


1
Expert's answer
2021-02-22T17:21:16-0500

Price =460,000

Quantity =90,000 cars

Price Elasticity =1.7(showing that the product is elastic)

New Price =415,000


Reduction is Price ="460,000-415,000=45,000"


Percentage reduction in Price ="\frac{45,000}{460,000}×100=9.78\%"


Percentage reduction in Price =9.78%


"Price Elasticity =\frac{\% \Delta Quantity} {\%\Delta Price}"


"=1.7=\frac{\% \Delta Quantity} {9.78}"


% change in quantity =1.7×9.78=16.63%


Thus the quantity demanded for the cars increased by 16.63%;


If 100%=90,000 cars

116.63%=?


="\frac{116.63×90,000}{100}" =104,967 cars.


Therefore, the annual sales will increase from 90,000 cars to 104,967cars since the average price declined, leading to increase in the quantity of cars demanded.